OpenStudy (anonymous):

Find the area enclosed by the curves y=7-2x^2 and y=x^2+4

OpenStudy (anonymous):

TO FIND THE AREA ENCLOSED YOU SHOULD SOLVE FOR COMMON POINTS y=7-2x^2 and y=x^2+4 7-2X^2=X^2+4 FIND THE VALUE OF X AND INTEGRATE BETWEEN THE TWO VALUES OF X FIND THE AREA USING THE BELOW FORMULA \[\int\limits_{X1}^{X2}F(Y1)-(F(Y2) DX\]

OpenStudy (anonymous):

thanks